# Conjugacy Class Growth in Virtually Abelian Groups

### v2

@@ Abstract.
Key words and phrases:
Conjugacy class, Virtually abelian groups, Affine Coxeter groups, Growth rates
2020 Mathematics Subject Classification:
05E16, 20E45, 20F55, 20F69
The first author was funded by Heilbronn Institute for Mathematical Research
The second author was funded by Heilbronn Institute for Mathematical Research
@@ 1. Introduction
Given a finitely generated group $G$ generated by some set $S$, every element in $G$ can be represented by some word over the alphabet $S\cup S^{-1}$. For an element $g\in G$, we define the length $\ell_{S}(g)$ of $g$ to be the minimal length of all words which represent $g$ with respect to the generating set $S$. This naturally leads to the study of the growth function of $G$, which counts the number of elements of length at most $n$. The asymptotic properties of growth functions of groups have been extensively studied with a highlight being Gromov’s celebrated theorem [7] characterising groups of polynomial growth as the virtually nilpotent groups.
The growth function can be generalised by restricting our attention to subsets of interest. In particular, taking some subset $U$ of $G$, we let $\beta_{U,S}(n)$ be the number of elements in $U$ whose length is at most $n$ and look at how this number changes as $n$ increases. The growth of subgroups of groups has been studied by a number of authors [3, 5, 12] but much less attention has been paid to other interesting subsets.
@@
Theorem 3.7.
@@
We emphasise that this is not to be confused with the conjugacy growth function, which counts the number of conjugacy classes intersecting the ball of radius $n$. The conjugacy growth function in virtually abelian groups was studied in [6]. Combinatorial properties including growth of other subsets of virtually abelian groups are explored in [4], and as noted in Remark 3.8, Theorem 3.7 also follows from general results in that article. However, the method of proof we use here is much more direct, and as we will see, allows for the degree of polynomial growth to be explicitly calculated.
To the authors’ knowledge, the first appearance of the growth of a conjugacy class is in [11], where it appears as a special case of the ‘orbital counting problem in conjugacy classes’. The authors of that paper are interested in examples of groups where all non-trivial conjugacy classes have equivalent growth functions, as well as the exponential growth rate of the function. They show that in the cases of hyperbolic groups, and the integer Heisenberg group, all infinite conjugacy classes have the same growth. In the present paper we show that this is not the case in general for virtually abelian groups (see Proposition 3.9), while providing an example of where it does hold (Example 3.10).
@@
Theorem 4.10.
@@
This article is organised in the following way. We start by giving preliminary definitions and basic results in group theory and growth in Section 2. Section 3 is devoted to showing the first of our two main results: (1) that the growth function of conjugacy classes in virtually abelian groups is polynomial. In Section 4, we give background definitions and results on affine Coxeter groups. The affine Coxeter groups are a particularly nice class of virtually abelian groups as they have a rich geometric structure to them. We end this section by showing the second of our two main results: (2) that the growth function of the conjugacy class of an element in an affine Coxeter group is polynomial with degree equal to the reflection length of its finite part.
@@ 2. Definitions and basic results
We begin by recalling some standard definitions in group theory and growth functions. The interested reader is referred to [10] for a more thorough introduction to the growth of groups.
@@ 2.1. Growth in groups
We start by giving some basic notions on words representing elements of a group and the growth rate of a set in relation to a group.
Definition 2.1.
Let $G$ be a finitely generated group and $S$ a choice of finite generating set. We let $\left\{S\cup S^{-1}\right\}^{\star}$ denote the words over the alphabet $S$ and its inverses, including the empty word. Given a word $w\in\left\{S\cup S^{-1}\right\}^{\star}$ we let $\lvert w\rvert$ denote the length of the word $w$ and we write $w=_{G}g$ if $w$ represents $g$ in the group $G$. Many words can represent the same element, but there will always be some word of minimal length representing a given $g$. The word length of an element $g\in G$ with respect to $S$ is then defined to be this minimal length
@@
Proof.
For elements $h,x,y\in G$ we have $d_{G,S}(x,y)=\ell_{S}(x^{-1}y)=\ell_{S}(x^{-1}h^{-1}hy)=d_{G,S}(hx,hy)$, and so the word metric is invariant under left multiplication. In other words $d_{G,S}(1,x)\leq n$ if and only if $d_{G,S}(h,hx)\leq n$ and we have a length-preserving bijection between the elements of the $n$-ball centred at $1$ and the elements of the $n$-ball centred at $h$. ∎
@@
for all $n\in{\mathbb{N}}$. If $f\preccurlyeq g$ and $g\preccurlyeq f$ then we write $f\sim g$ and say that the functions are equivalent. Note that this defines an equivalence relation. We write $f\prec g$ if $f\preccurlyeq g$ but it is not the case that $f\sim g$.

### v3

@@ Abstract.
2020 Mathematics Subject Classification. 05E16, 20E45, 20F55, 20F69.
@@ 1. Introduction
Given a finitely generated group $G$ generated by some set $S$, every element in $G$ can be represented by some word over the alphabet $S\cup S^{-1}$. For an element $g\in G$, we define the length $\ell_{S}(g)$ of $g$ to be the minimal length of all words which represent $g$ with respect to the generating set $S$. This naturally leads to the study of the growth function of $G$, which counts the number of elements of length at most $n$. The asymptotic properties of growth functions of groups have been extensively studied with a highlight being Gromov’s celebrated theorem [Gromov] characterising groups of polynomial growth as the virtually nilpotent groups.
The growth function can be generalised by restricting our attention to subsets of interest. In particular, taking some subset $U$ of $G$, we let $\beta_{U,S}(n)$ be the number of elements in $U$ whose length is at most $n$ and look at how this number changes as $n$ increases. The growth of subgroups of groups has been studied by a number of authors [CantrellSharp, DavisOlshanskii, Schesler] but much less attention has been paid to other interesting subsets.
@@
Theorem LABEL:thm:conj_class_poly.
@@
We emphasise that this is not to be confused with the conjugacy growth function, which counts the number of conjugacy classes intersecting the ball of radius $n$. The conjugacy growth function in virtually abelian groups was studied in [Evetts]. Combinatorial properties including growth of other subsets of virtually abelian groups are explored in [RationalVAb], and as noted in Remark LABEL:rem:altproof, Theorem LABEL:thm:conj_class_poly also follows from general results in that article. However, the method of proof we use here is much more direct, and as we will see, allows for the degree of polynomial growth to be explicitly calculated.
To the authors’ knowledge, the first appearance of the growth of a conjugacy class is in [ParkPaul], where it appears as a special case of the ‘orbital counting problem in conjugacy classes’. The authors of that paper are interested in examples of groups where all non-trivial conjugacy classes have equivalent growth functions, as well as the exponential growth rate of the function. They show that in the cases of hyperbolic groups, and the integer Heisenberg group, all infinite conjugacy classes have the same growth. In the present paper we show that this is not the case in general for virtually abelian groups (see Proposition LABEL:prop:allgrowths), while providing an example of where it does hold (Example LABEL:ex:kleinbottle).
@@
Theorem LABEL:thm:poly_growth_Coxeter.
@@
This article is organised in the following way. We start by giving preliminary definitions and basic results in group theory and growth in Section 2. In Section LABEL:sec:coxeter_groups, we give background definitions and results on affine Coxeter groups. The affine Coxeter groups are a particularly nice class of virtually abelian groups as they have a rich geometric structure to them. Section LABEL:sec:virtually_abelian_groups is then devoted to showing our two main results: (1) that the growth function of conjugacy classes in virtually abelian groups is polynomial and (2) that the growth function of the conjugacy class of an element in an affine Coxeter group is polynomial with degree equal to the reflection length of its finite part.
@@ 2. Definitions and basic results
We begin by recalling some standard definitions in group theory and growth functions. The interested reader is referred to [Loh] and [Mann] for a more thorough introduction to the growth of groups.
@@ 2.1. Growth in groups
We start by giving some basic notions on words representing elements of a group and the growth rate of a set in relation to a group. {defi} Let $G$ be a finitely generated group and $S$ a choice of finite generating set. We let $\left\{S\cup S^{-1}\right\}^{\star}$ denote the words over the alphabet $S$ and its inverses, including the empty word. Given a word $w\in\left\{S\cup S^{-1}\right\}^{\star}$ we let $\lvert w\rvert$ denote the length of the word $w$ and we write $w=_{G}g$ if $w$ represents $g$ in the group $G$. Many words can represent the same element, but there will always be some word of minimal length representing a given $g$. The word length of an element $g\in G$ with respect to $S$ is then defined to be this minimal length
@@
Proof 2.1.
For $h,x,y\in G$ we have $d_{G,S}(x,y)=\ell_{S}(x^{-1}y)=\ell_{S}(x^{-1}h^{-1}hy)=d_{G,S}(hx,hy)$, and so the word metric is invariant under left multiplication. In other words $d_{G,S}(1,x)\leq n$ if and only if $d_{G,S}(h,hx)\leq n$ and we have a length-preserving bijection between the elements of the $n$-ball centred at $1$ and the elements of the $n$-ball centred at $h$.
@@
for all $n\in{\mathbb{N}}$. If $f\preccurlyeq g$ and $g\preccurlyeq f$ then we write $f\sim g$ and say that the functions are
Definition 2.
equivalent. Note that this defines an equivalence relation. We write $f\prec g$ if $f\preccurlyeq g$ but it is not the case that $f\sim g$.